Solve system of eqautions by subsitution
Kirsten is five years older then Tom. The product of their ages is 24. How old are Kirsten and Tom? make your system of equations, then solve.

Found 2 solutions by TutorGeek, Alan3354:
Answer by TutorGeek(1)

Well i solved this mentally and what i got is that Kristen is 8 and Tom is 3 because Kristen is 5 years older than Tom and 3+5=8...the product of their ages is supposed to be 24 and 8*3=24. ^_^

Answer by Alan3354(69443)

You can make 2 equations, then sub, etc, then you get a quadratic.
To factor it, you find a pair of factors of 24 that differ by 5, so why not just do that?
3*8 = 24